OpenStudy (amtran_bus):

Find the arc length of ellipse x=4sintheta and y=3costheta

OpenStudy (amtran_bus):

So I know you can say\[\int\limits_{a}^{b}\sqrt{1+(dy/dx)^2}\]

OpenStudy (amtran_bus):

and also \[\frac{ x }{ 4 }=\sin \theta , \frac{ y }{ 3 }= \cos \theta\]

OpenStudy (amtran_bus):

so (x/4)^2 +(y/3)^2 =1

OpenStudy (amtran_bus):

Taking the derivative, you got 4 cos theta and -3 sin theta

OpenStudy (amtran_bus):

@amistre64

OpenStudy (amistre64):

id use: \[\int \sqrt{(x')^2+(y')^2}~dt\]

OpenStudy (amtran_bus):

That isnt the formula in my book? How did you get it and why does it work? Just from the trig identity?

OpenStudy (amistre64):

its the general setup for the arclength "s", which is just the pythag thrm s, the section, squared; is equal to sum of the square of the parts  s^2 = x^2 + y^2, or simply s = sqrt(x^2 + y^2) as s and x and y get small ...

OpenStudy (amistre64):